OpenStudy (anonymous):

15n + 16 + 4n = 130 x= 19.12, right?

OpenStudy (anonymous):

the answer should be 10.36 (rounded)

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

the answer is 6

OpenStudy (anonymous):

..... *confused*

OpenStudy (anonymous):

15n + 4n = 130 - 16 19n = 114 n = 114/19 n = 6

jimthompson5910 (jim_thompson5910):

15n + 16 + 4n = 130 19n + 16 = 130 19n + 16-16 = 130-16 ... Subtract 16 from both sides. 19n + 0 = 114 19n = 114 19n/19 = 114/19 ... Divide both sides by 19 to isolate n. 1n/1 = 6 1n = 6 n = 6

OpenStudy (anonymous):